OpenStudy (anonymous):

**MEDAL FOR BEST ANSWER!!** @mathstudent55

OpenStudy (anonymous):

Factor this expression completely, then place the factors in the proper location on the grid. x^6 - y^6

OpenStudy (mathstudent55):

Notice that the exponents are both 6. Is x^6 the cube of something? Is y^6 the cube of something?

OpenStudy (anonymous):

I have no idea what it is the cube of.

OpenStudy (mathstudent55):

Remember this rule: \((a^m)^n = a^{mn} \) Example: \((x^4)^3 = x^{12} \)

OpenStudy (anonymous):

So it's the cube of 2?

OpenStudy (mathstudent55):

Now think of \(x^6\). When you raise a power to a power, you multiply powers, so since 6, the exponent , is equal to 2 * 3, \(x^6 = (x^2)^3\).

OpenStudy (mathstudent55):

That means x^6 is the cube of x^2.

OpenStudy (anonymous):

Okay so x^2-y^2. . . ?

OpenStudy (mathstudent55):

You are on the right track, but I need to change a little the steps. Let's factor it it this way first. Just like \(x^6 = (x^2)^3\) \(x^6 = (x^3)^2 \), right?

OpenStudy (anonymous):

Yeah. So how do I put that in the equation? I have a graph that looks like: _x_____-_)_x________________^2_ and I need to figure out what goes in the blanks (which are the _)

OpenStudy (mathstudent55):

Let's rewrite x^6 as the square of something instead of as the cube of something.

OpenStudy (mathstudent55):

\(x^6 - y^6 = (x^3)^2 - (y^3)^2 \) Now let's factor the difference of two squares first. \( (x^3)^2 - (y^3)^2 \) \(= (x^3 + y^3)(x^3 - y^3) \)

OpenStudy (mathstudent55):

Now we need to factor the sum of cubes and the difference of cubes.

OpenStudy (anonymous):

How do I do that?

OpenStudy (mathstudent55):

\(= (x^3 + y^3)(x^3 - y^3)\) \( (x + y)(x^2 - xy + y^2)(x - y)(x ^2 + xy + y^2 )\)

OpenStudy (anonymous):

That doesn't fit in my spaces though. It's _x_____-_)_x________________^2_

OpenStudy (mathstudent55):

That's the complete factoring.

OpenStudy (anonymous):

Okay so how do I get to the complete factoring?

OpenStudy (mathstudent55):

What I did is the complete factoring of the expression x^6 - y^6. That's what the problem asked for.

OpenStudy (anonymous):

That doesnt fit in the spaces though.

OpenStudy (mathstudent55):

Since your earlier problem today was about factoring the difference of 2 cubes, maybe all this problem is looking for is for it to be factored as the difference of two cubes. If so, then you do this: \(x^6 - y^6 = (x^2)^3 - (y^2)^3 = (x^2 - y^2)(x^4 + x^2y^2 + y^4)\) The problem here is that it is not completely factored as the instructions ask to do.

OpenStudy (mathstudent55):

Here's an example of a difference of cubes: \(8y^3 - 27 \) \(= (2y)^3 - 3^3 \) \( =(2y - 3)[ (2y)^2 + 2y \cdot 3 + 3^2\)] \( = (2y - 3)(4y^2 + 6y + 9) \)

OpenStudy (anonymous):

Let me send you a picture. Hold on. @mathstudent55

OpenStudy (anonymous):

I need the answer to fit in those boxes.